### v1

# Composite Fermion Theory of Fractional Chern Insulator Stability

Abstract
We develop a mean-field theory of the stability of fractional Chern insulators based on the dipole picture of composite fermions (CFs). We construct CFs by binding vortices to Bloch electrons and derive a CF single-particle Hamiltonian that describes a Hofstadter problem in the enlarged CF Hilbert space, with the well-known trace condition emerging naturally in the small-$q$ limit. Applied to twisted MoTe2, the calculated CF phase diagram matches closely with that from exact diagonalization, and the projected many-body wavefunctions achieve exceptionally high overlaps with the latter. Our theory provides both a microscopic understanding and a computationally efficient tool for identifying fractional Chern insulators.
Introduction.—Fractional Chern insulators (FCIs) are lattice analogs of the fractional quantum Hall effect (FQHE), but realized in zero magnetic field. Proposed over a decade ago [1, 2, 3, 4, 5, 6], FCIs have only recently been observed experimentally in twisted MoTe2 bilayers [7, 8, 9, 10] and in pentalayer graphene on hBN [11]. Unlike conventional FQHE platforms, these moiré systems offer a wide range of tuning knobs—such as twist angle, gate voltage, and dielectric environment—enabling exploration of a much richer phase diagram [7, 8, 9, 10, 11, 12, 13]. Given this experimental progress, understanding the stability of FCIs has become an urgent issue.
To date, two main theoretical approaches have been pursued. Numerically, techniques such as exact diagonalization (ED) and density matrix renormalization group (DMRG) have been used to map out FCI phase diagrams, but these are computationally expensive and limited to small system sizes. Analytical approaches, in contrast, have primarily focused on the sufficient conditions to realize FCI phases, mostly by mimicking the lowest Landau level (LLL) physics [14]. Notable examples include the trace condition $\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}=|\Omega_{\bm{k}}|$ restoring the Girvin-MacDonald-Platzman (GMP) algebra [15] in the small-$q$ limit [16, 17, 18, 19, 20], the ideal band condition yielding $\bm{k}$-holomorphic Bloch wavefunctions [21, 22, 23], and the notion of vortexability extending the flux attachment picture in combination of band projection [24, 25]. While insightful, these criteria apply only to narrow regions of parameter space similar to the LLL. In fact, recent numerical studies have shown that deviating from such conditions can, in some cases, enhance the many-body gap [26, 27]. This motivates a central question: Can we develop a physically transparent and computationally efficient framework for diagnosing FCI stability beyond the ideal limit?
In this Letter, we present a quantitative theory of FCI stability based on the dipole picture of composite fermions (CFs) [28, 29, 30, 31, 32, 33, 34]. We construct CFs on a lattice by binding Bloch electrons to vortices residing in their own LLL, and derive an effective CF Hamiltonian using the method of preferred-charge substitution [31, 35, 36]. In the small-$q$ limit, the resulting Hamiltonian describes a Hofstadter problem in the enlarged CF Hilbert space, where both the scalar potential and the magnetic field are periodic (see also Ref. [34]). In particular, assuming a flat electron band, the trace condition further reduces this Hamiltonian to the Aharonov-Casher form [37], which guarantees the exact flatness of the lowest CF band. For general cases, we use the GMP expansion [32, 38], combined with our recently developed lattice projective construction of FCI states [33], to extract CF band structures. CF phase diagrams can then be obtained by tracking the closing of the CF band gap.
As a concrete example, we apply our theory to study FCI stability in twisted MoTe2. Using only the CF single-particle Hamiltonian, we obtain a phase diagram that captures both the shape of the FCI region and the location of the maximal many-body gap, in agreement with ED results. We show that the dependence on Coulomb interaction strength and twist angle can be understood by examining the evolution of the CF band structure. We further validate our approach by constructing many-body electron wavefunctions through projection of the CF states, and find excellent overlaps with ED ground states. Our CF-based framework thus provides an efficient and analytically tractable tool for diagnosing FCI stability, offering predictive power and physical insight into the interplay between band geometry and interactions.
The Preferred CF Hamiltonian.—Although not as widely known as the flux attachment formalism [39, 40], the dipole picture offers a direct and physically motivated route to constructing the CF Hamiltonian. In the context of the FQHE, a CF dipole is formed by binding each electron to a $2s$-fold vortex [28]. Each vortex carries a charge opposite to that of the electron, $q_{v}=-c^{2}q_{e}$, where $c^{2}\equiv 2s\nu$ and $\nu$ is the electron filling factor. As a result, the total CF charge becomes $q_{\text{CF}}=(1-c^{2})q_{e}$. When $\nu$ lies in the Jain sequence $\nu=p/(2ps+1)$, the CF filling factor becomes an integer $p$, and the FQHE of electrons maps to the integer quantum Hall effect of CFs. This construction underlies the extended Hamiltonian theory of Murthy and Shankar, which provides a powerful tool for studying the FQHE [31]. In what follows, we apply it to FCIs and present a self-contained derivation within this framework; readers may consult Ref. [31, 33] for more background.
Our starting point is the electronic Hamiltonian projected into a single band (we omit the band index for simplicity):
$H=H_{0}+U=\sum_{\bm{k}}\varepsilon_{\bm{k}}c_{\bm{k}}^{\dagger}c_{\bm{k}}+\frac{1}{2}\sum_{\bm{q}}V_{\bm{q}}:\bar{\rho}_{e,\bm{q}}\bar{\rho}_{e,-\bm{q}}:\;,$ (1)
where $\varepsilon_{\bm{k}}$ is the band dispersion, $V_{\bm{q}}$ is the Coulomb interaction, and $\bar{\rho}_{e,\bm{q}}=\sum_{i=1}^{N}P_{e}e^{i\bm{q}\cdot\hat{\bm{r}}_{e,i}}P_{e}$ is the projected electron density operator, with $|P_{e}=\sum_{\bm{k}}\ket{\psi_{\bm{k}}}\bra{\psi_{\bm{k}}}$ the Bloch band projector. We assume that the band carries Chern number C=1C=-1 and the electron filling factor falls within the Jain sequence $\nu=p/(2sp+1)$.
To construct CFs on a lattice, we introduce vortices as auxiliary degrees of freedom similar to the FQHE. Although we continue to call them “vortices”, they are best viewed as fictitious particles that carry fractional charge $q_{v}=-c^{2}q_{e}$, and experience an effective magnetic field set by the average Berry curvature flux, $B_{\text{eff}}=h/(eA_{\text{u.c.}})$, where $A_{\text{u.c.}}$ is the unit cell area. The vortex unit cell, determined by magnetic translation symmetry, is thus enlarged by a factor of $(2ps+1)$ relative to the electron unit cell.
The CF is a bound state of an electron and a vortex that lives in the enlarged single-particle Hilbert space
$\mathcal{H}_{\text{CF}}=\mathcal{H}_{e}\otimes\mathcal{H}_{v}\;,$ (2)
where $\mathcal{H}_{e}$ is spanned by the Bloch states and $\mathcal{H}_{v}$ by the vortex LLL states. The “binding” between a Bloch electron and a vortex is achieved by replacing the electron density operator with the so-called preferred charge density operator [31, 35, 36], reflecting the dipole structure of CFs:
$\bar{\rho}_{e,\bm{q}}\mapsto\bar{\rho}^{P}_{\bm{q}}=\bar{\rho}_{e,\bm{q}}\otimes\mathbb{1}_{v}-c^{2}\mathbb{1}_{e}\otimes\bar{\rho}_{v,-\bm{q}}\;,$ (3)
where $\bar{\rho}_{v,\bm{q}}$ is the LLL-projected vortex density operator:
$\bar{\rho}_{v,\bm{q}}=\sum_{i=1}^{N}P_{v}e^{-i\bm{q}\cdot\hat{\bm{r}}_{v,i}}P_{v}=e^{-(q\ell_{v})^{2}/4}\sum_{i=1}^{N}e^{-i\bm{q}\cdot\hat{\bm{R}}_{v,i}}\;,$ (4)
with the vortex guiding center coordinates $\hat{\bm{R}}_{v}=P_{v}\hat{\bm{r}}_{v}P_{v}$, and the corresponding magnetic length $\ell_{v}=\sqrt{\hbar/|q_{v}B_{\text{eff}}|}$. The minus sign in the exponent of $\bar{\rho}_{v,\bm{q}}$ reflects the fact that the magnetic translation of a negatively charged vortex is equivalent to a reversed translation of the positive charge.
To carry out the preferred charge substitution, we first unpack the normal-ordered electron interaction in Eq. 1 and replace $\bar{\rho}_{e,\bm{q}}$ with $\bar{\rho}^{P}_{\bm{q}}$. We then re-normal order the resulting Hamiltonian in terms of $\bar{\rho}^{P}_{\bm{q}}$ to recover the standard form of an interacting Hamiltonian, arriving at (see Supplementary Material (SM) [41] for details):
$H^{P}=H^{P}_{\text{single}}+H^{P}_{\text{int}}$ $=H_{0}+H_{0}^{P}+H^{P}_{\text{int}}\;,$ (5)
where $H_{0}$ is the electron band dispersion, $H^{P}_{\text{int}}=\sum_{\bm{q}}V_{\bm{q}}:\bar{\rho}^{P}_{\bm{q}}\bar{\rho}^{P}_{-\bm{q}}:$ is the normal-ordered CF interaction, and $H_{0}^{P}$ is a new single-particle term arising from the normal-ordering:
$H_{0}^{P}=\sum_{\bm{k},\bm{k}^{\prime}}\sum_{z_{v},z_{v}^{\prime}}H_{0}^{P}(\bm{k},z_{v};\bm{k}^{\prime},z_{v}^{\prime})\,c_{\bm{k},z_{v}}^{\dagger}c_{\bm{k}^{\prime},z_{v}^{\prime}}\;.$ (6)
[…]
$H_{0}^{P}(\bm{k},z_{v};\bm{k}^{\prime},z_{v}^{\prime})=\frac{-c^{2}}{2}\sum_{\bm{q}}V_{\bm{q}}\langle\bar{\varrho}_{e,\bm{q}}\otimes\bar{\varrho}_{v,\bm{q}}+\text{h.c.}\rangle\;,$ (7)
where $\bar{\varrho}_{\alpha,\bm{q}}=e^{i\mathop{\mathrm{sgn}}(\alpha)\bm{q}\cdot\hat{\bm{r}}_{\alpha}}$ denotes the matrix element of the projected density operator for species $\alpha=e,v$ with $\mathop{\mathrm{sgn}}(e)=1$ and $\mathop{\mathrm{sgn}}(v)=-1$.
Equations 5, 6 and 7 constitute the core result underlying our CF theory. They describe the band structure of CFs and provide the foundation for analyzing FCI stability. Crucially, at filling $\nu=p/(2ps+1)$, the CFs occupy exactly $p$ bands, mapping the problem of an FCI of electrons to that of an integer Chern insulator of CFs. This follows from the requirement that the CF periodicity must be compatible with both the electron and vortex unit cells, leading to an enlarged unit cell and hence $p$ filled CF bands.
In the remainder of this Letter, we demonstrate that much of the CF physics is already captured by the CF single-particle Hamiltonian $H^{P}_{\text{single}}=H_{0}+H_{0}^{P}$, which encodes interaction effects through its dependence on the Coulomb potential $V_{\bm{q}}$.
To diagonalize $H^{P}_{\text{single}}$, we can expand $\bar{\rho}_{e,\bm{q}}$ in terms of GMP density operators [32]:
$\bar{\rho}_{e,\bm{q}}=\sum_{\bm{G}}c(\bm{q},\bm{G})\rho^{\text{GMP}}_{e,[\bm{q}]+\bm{G}}\;,$ (8)
where $\bm{G}$ is the electron reciprocal lattice vector, $\bm{q}]$ denotes $\bm{q}$ folded back to the first Brillouin zone (BZ), and the coefficients
$c(\bm{q},\bm{G})=\frac{1}{N_{\phi,e}}\mathop{\mathrm{Tr}}\left[\bar{\varrho}_{e,\bm{q}}\varrho^{\text{GMP}}_{e,-[\bm{q}]-\bm{G}}\right]$ (9)
encode all lattice effects specific to the FCI. This so-called GMP expansion is complete and can be used for any fermionic bilinear [32, 38]. The resulting CF band structure problem over a finite-size sample on a torus can then be solved using the lattice projective construction as described in Ref. [33].
Small-$q$ Limit and the Trace Condition.—It is instructive to first consider the situation when the form factor of the density operator decays exponentially, or when $V_{\bm{q}}$ is appreciable only at small $\bm{q}$. For instance, the vortex density operator $\bar{\rho}_{v,\bm{q}}$ carries a form factor $e^{-(q\ell_{v})^{2}/4}$ from the LLL projection [see Eq. 4]. We can then expand the projected density operators for both species $\alpha=e,v$ to second order in $q$:
$\bar{\varrho}_{\alpha,\bm{q}}\simeq P_{\alpha}+i\,\mathop{\mathrm{sgn}}(\alpha)\,q_{\mu}P_{\alpha}\hat{r}_{\alpha}^{\mu}P_{\alpha}-\frac{1}{2}q_{\mu}q_{\nu}P_{\alpha}\hat{r}_{\alpha}^{\mu}\hat{r}_{\alpha}^{\nu}P_{\alpha}\;.$ (10)
Assuming further that $V_{\bm{q}}$ is isotropic, then only terms $\sim\bm{q}^{2}$ survive after integrating over $\bm{q}$. Up to a constant, the CF single-particle Hamiltonian takes the form (see SM [41]):
$\begin{split}H^{P}_{\text{single}}(\bm{k},z_{v})&=\varepsilon_{\bm{k}}+\frac{\hbar^{2}}{2m_{*}\ell_{e}^{2}\ell_{v}^{2}}(\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}+\Omega_{\bm{k}})\\ &\quad+\frac{\hbar^{2}}{2m_{*}\ell_{e}^{2}\ell_{v}^{2}}(2\mathcal{D}-z_{v})(2\bar{\mathcal{D}}-\hat{\bar{z}}_{v})\;,\end{split}$ (11)
where $\ell_{\alpha}=\sqrt{\hbar/|q_{\alpha}B_{\text{eff}}|}$ and the CF effective mass is generated entirely by the Coulomb interaction:
$\frac{\hbar^{2}}{2m_{*}\ell_{e}^{2}\ell_{v}^{2}}=\frac{c^{2}}{4}\sum_{\bm{q}}V_{\bm{q}}\bm{q}^{2}e^{-\lambda^{2}q^{2}}\;.$ (12)
We have added a $\lambda$-parameterized exponential regulator to ensure convergence (see SM [41] for its origin).
The terms proportional to $1/m_{*}$ in Eq. (11) reflect geometric properties of the underlying electron band. The complex covariant derivative is defined as $\mathcal{D}=\tfrac{1}{2}(\mathcal{D}_{x}-i\mathcal{D}_{y})$, where $\mathcal{D}_{\mu}=i\partial_{k_{\mu}}+A_{\mu}(\bm{k})$, and $A_{\mu}(\bm{k})=i\braket{u_{\bm{k}}|\partial_{k_{\mu}}u_{\bm{k}}}$ is the Berry connection. The operator $\hat{\bar{z}}_{v}=-2\ell_{v}^{2}\partial_{z}$ is defined so that the vortex guiding-center coordinates satisfy $\hat{R}_{v,x},\hat{R}_{v,y}]=i\ell_{v}^{2}$. Interestingly, the trace condition expression $\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}+\Omega_{\bm{k}}$, where $\mathcal{G}_{\bm{k},\mu\nu}=\mathop{\mathsf{Re}}[\braket{\partial_{k_{\mu}}u_{\bm{k}}|\partial_{k_{\nu}}u_{\bm{k}}}]-A_{\mu}A_{\nu}$ is the quantum metric and $\Omega_{\bm{k}}=\mathop{\mathsf{Im}}[\braket{\partial_{k_{x}}u_{\bm{k}}|\partial_{k_{y}}u_{\bm{k}}}]$ is the Berry curvature, appears naturally in the CF Hamiltonian. Had we assumed the electron band carries positive Chern number $C=1$, the sign of $\Omega_{\bm{k}}$ would be reversed.
Taken together, Eq. 11 describes a Hofstadter problem in the $(\bm{k},z_{v})$ space: the first two terms act as a periodic scalar potential, while the last term, taking the form of an Aharonov-Casher (AC) Hamiltonian [37], describes a periodic magnetic field. It is clear that if the electron band is flat ($\varepsilon_{\bm{k}}=0$) and the so-called trace condition $\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}+\Omega_{\bm{k}}=0$ is satisfied, then the Hamiltonian reduces to a pure AC form, which guarantees that the lowest CF band is exactly flat. This perspective yields several new insights. First, contrary to earlier studies where $\varepsilon_{\bm{k}}$ was neglected, a nonzero electron band dispersion can help stabilize the FCI phase by compensating for deviations from the trace condition. Second, the trace condition is only meaningful when the lowest CF band is filled; when multiple CF bands are filled, it no longer applies as a stability criterion.
We note that a Hamiltonian of the same form as Eq. 11 was proposed in Ref. [34], based on fitting a postulated quadratic electron-vortex dipole model. Our result provides a first-principle derivation of this structure within the preferred charge formalism.
Figure 1: CF phase diagram (a), $6\times 4$ ED phase diagram (b), geometric fluctuations (c), and typical CF bands in twisted MoTe2 (d–f). The color maps in (a) and (b) represent the CF band gap and the many-body gap obtained from ED, respectively. Panel (c) shows the standard deviation of the effective magnetic field $\Omega_{\bm{k}}$ and the effective scalar potential of the CF Hamiltonian (11). The standard deviation of the trace condition is also shown. Panels (d)–(f) show the two lowest CF bands for three representative points labeled in (a), computed on a $9\times 9$ electronic sample, with colors represents the band widths. Note: For visualization, the CF bands are unfolded into the electronic Brillouin zone (BZ), rather than plotted within the CF’s original BZ, which occupies only one-third of the electronic BZ at filling $\nu=1/3$. The extra threefold degeneracy within CF’s reduced BZ, known as the symmetry fractionalization pattern due to the interplay between topological orders and magnetic translation symmetries [42, 43, 44, 33], is also explicitly visible in our calculation.
CF Band Structure and CF Phase Diagram in tMoTe2.—We now apply our theory to the continuum model of the $R$-stacked twisted MoTe2 bilayer. We assume the system is valley polarized, and focus on the effective Hamiltonian at the $K$ valley, given by [45]:
$H_{K}=\begin{bmatrix}h_{b}(\bm{r})&\Delta_{T}(\bm{r})\\ \Delta_{T}^{\dagger}(\bm{r})&h_{t}(\bm{r})\end{bmatrix}$, (13)
where $h_{b,t}=-(\bm{k}-\bm{K}_{b,t})^{2}/2m_{\text{eff}}+2v\sum_{i=1,3,5}\cos(\bm{g}_{i}\cdot\bm{r}\pm\psi)$ is the top/bottom layer Hamiltonian with a moiré potential, $\Delta_{T}=w(1+e^{-i\bm{g}_{2}\cdot\bm{r}}+e^{-i\bm{g}_{3}\cdot\bm{r}})$ is the interlayer tunneling, and $\bm{g}_{i}=\frac{4\pi}{\sqrt{3}a/\theta}(\cos\frac{\pi(i-1)}{3},\sin\frac{\pi(i-1)}{3})$ are the moiré reciprocal vectors. The momenta $\bm{K}_{b}=\frac{1}{3}(\bm{g}_{4}+\bm{g}_{5})$ and $\bm{K}_{t}=\frac{1}{3}(\bm{g}_{3}+\bm{g}_{4})$ are the $K$ points of the top/bottom layer and $\theta$ is the twist angle. All parameters are from Ref. [46]. The Coulomb potential with dual-gate screening is taken to be $V_{\bm{q}}=\frac{1}{A_{\text{u.c.}}}\frac{e^{2}\tanh(|\bm{q}|d)}{2\epsilon\epsilon_{0}|\bm{q}|}$, where $d$ is the gate-to-sample distance.
Let us focus on the hole filling $\nu=-2/3$ of the topmost valence band. We take a particle-hole transformation to convert it to an equivalent electron-doped model at filling $\nu=1/3$, squeezing out a Hartree-term $\varepsilon_{\bm{k}}^{\text{Hartree}}\equiv-\sum_{\bm{G}}V_{\bm{G}}(\sum_{\bm{k}^{\prime}}\bar{\varrho}_{-\bm{G}}([\bm{k}^{\prime}]))\bar{\varrho}_{\bm{G}}([\bm{k}])$ that is absorbed into the original band dispersion $\varepsilon_{\bm{k}}\mapsto-\varepsilon_{\bm{k}}-\varepsilon_{\bm{k}}^{\text{Hartree}}$. We construct the CF Hamiltonian using the preferred charge substitution (3), and solve it via the GMP expansion (8) and our recently developed lattice projective construction [33].
Figure 1(d) shows the CF band structure at $\theta=3.55^{\circ}$ and $\epsilon=15$. Only the two lowest CF bands are shown, since at $\nu=1/3$ only the lowest band is fully filled, with all higher bands empty. We observe that both bands are flat, reflecting the near-ideal geometric conditions of the underlying Bloch band: the trace condition is nearly saturated, and the fluctuations in Berry curvature is small [see Fig. 1(c)]. The electron band dispersion is also small. In this case, the CF Hamiltonian Eq. 11 describes motion of a charged particle in a uniform magnetic field in the enlarged ($\bm{k},z_{v})$ space, and the CF bands closely resemble the flat CF Landau levels in the FQHE.
In the non-interacting CF limit, the CF band gap provides a direct estimate of the activation energy—the cost of creating a well-separated quasiparticle–quasihole pair. We plot this gap as a function of twist angle $\theta$ and dielectric constant $\epsilon$ in Fig. 1(a). The resulting CF phase diagram shows remarkable agreement with ED results in Fig. 1(b), capturing not only the shape of the FCI region but also the location of the maximal many-body gap. This agreement strongly supports the predictive power of our CF-based approach and confirms that the CF band gap serves as a reliable proxy for FCI stability.
To understand the breakdown of the FCI phase, we track the evolution of the CF band structure along the vertical line in Fig. 1(a), varying $\epsilon$ at fixed $\theta=3.55^{\circ}$. In FQHE, neglecting Landau level mixing, the many-body gap scales with the interaction strength and remains open until the interaction is completely turned off. FCIs, however, behaves differently. As $\epsilon$ increases, the inverse CF mass $1/m_{*}$ decreases [see Eq. 12], leading to two key effects illustrated in Fig. 1(e). First, the second CF band shifts downward, reflecting the expected reduction in interaction energy. Second, the weakened interaction amplifies the influence of the electron band dispersion $\varepsilon_{\bm{k}}$, causing the CF bands to become increasingly dispersive. Once $\epsilon$ exceeds a critical value, the top and bottom CF bands touch and the gap closes, signaling the breakdown of the FCI phase. Notably, this transition cannot be captured by geometric criteria such as ideal-band condition or vortexability, which remain unchanged along this trajectory. In contrast, the electron band dispersion enters naturally into the CF Hamiltonian, allowing our framework to account for this interaction-driven destabilization.
Next, we fix the dielectric constant and vary the twist angle $\theta$. As shown in Fig. 1(c), both the scalar potential [the first two terms in Eq. 11] and the effective magnetic field $\Omega_{\bm{k}}$ exhibit increasing fluctuations with larger $\theta$. As a result, CF bands develop significant dispersion, as shown in Fig. 1(f) for $\theta=4.5^{\circ}$. In contrast, the band center shifts only slightly, reflecting the fact that the effective mass remains approximately constant. Thus, the enhanced fluctuations in the CF band energy are the primary cause of the decreasing CF band gap at larger twist angles.
[…]
Physical Wavefunction from Projection.—Our CF theory is formulated in an enlarged Hilbert space that includes auxiliary vortex degrees of freedom. To obtain the physical electronic wavefunction, we must project out these unphysical degrees of freedom. This projection is implemented as [34, 33]:
[…]
As shown in Fig. 2(b), at filling $\nu=1/3$, there are three nearly degenerate ground states characterized by evenly spaced center-of-mass momenta [49, 50]. We generate the projected electronic states at these momenta and plot their overlap with the ED ground states in Fig. 2(a). The agreement is remarkable: within the FCI phase, the projected states achieve overlaps exceeding 90%, with a peak value of $(99.37\pm 0.06)\%$ at $\theta=3.5^{\circ}$, where the many-body gap reach its maximum. As $\theta$ increases, the energy of eigenstates outside the topological manifold descend and eventually become the new ground states, signaling the breakdown of the FCI phase [Fig. 2(c)]. We continue to track the states at the three center-of-mass momenta where FCI’s topological ground states reside under this transition, and find that their overlaps with the projected electronic states remain finite. In contrast, the overlaps with the new ground states are negligible. This result confirms that our framework not only qualitatively reproduces the FCI phase diagram, but also provides quantitatively accurate many-body wavefunctions through the electronic projection defined in Eq. 14.
Summary.—We have developed a mean-field theory of FCI phase stability based on the dipole picture of CFs. By introducing vortices as auxiliary degrees of freedom and implementing the preferred charge substitution, we derived a CF single-particle Hamiltonian that describes a Hofstadter problem in the enlarged CF Hilbert space. In the small-$q$ limit, the trace condition emerges naturally from this construction as the criterion for flatness of the lowest CF band. The calculated CF phase diagram closely matches that from ED, and the projected many-body wavefunctions achieve high overlap with ED results. This agreement strongly supports the validity of our microscopic framework. Moreover, the CF band structure is obtained by diagonalizing a single-particle Hamiltonian, making our approach both reliable and efficient for predicting FCI phases.
In this work, we have neglected the residual CF interactions $H^{P}_{\text{int}}$, in part because the preferred charge substitution already incorporates essential interaction effects—most notably, the CF effective mass is entirely Coulomb-driven. An important next step is to include $H^{P}_{\text{int}}$ to study collective excitations such as the magnetoroton mode. Another promising direction is to explore whether residual CF interactions can drive symmetry-breaking phases, such as CF charge orders or CF pairings, beyond the simple band picture. Additionally, incorporating band-mixing effects remains an important task. These extensions could further deepen our understanding of FCI physics and strengthen the connection between microscopic theory and experimental observations.
Acknowledgment.—We thank Junkai Dong, Zhihuan Dong, Manato Fujimoto, Junren Shi, Jingtian Shi, and Jie Wang for stimulating discussions, and Junren Shi for sharing unpublished result. This work is supported by the Center on Programmable Quantum Materials, an Energy Frontier Research Center funded by DOE BES under Award No. DE-SC0019443.

### v2

@@
 We develop a mean-field theory of the stability of fractional Chern insulators based on the dipole picture of composite fermions (CFs).
-We construct CFs by binding vortices to Bloch electrons and derive a CF single-particle Hamiltonian that describes a Hofstadter problem in the enlarged CF Hilbert space, with the well-known trace condition emerging naturally in the small-$q$ limit.
-Applied to twisted MoTe2, the calculated CF phase diagram matches closely with that from exact diagonalization, and the projected many-body wavefunctions achieve exceptionally high overlaps with the latter.
+We construct CFs by binding vortices to Bloch electrons and derive a CF single-particle Hamiltonian that describes a Hofstadter problem in the enlarged CF Hilbert space, with the trace-condition term emerging naturally in the small-$q$ limit as part of the CF Hamiltonian.
+Going beyond the small-$q$ limit, we apply our theory to twisted MoTe2 and calculate its CF band structures.
+The resulting CF phase diagram matches closely with that from exact diagonalization, and the projected many-body wavefunctions achieve exceptionally high overlaps with the latter.
 Our theory provides both a microscopic understanding and a computationally efficient tool for identifying fractional Chern insulators.
@@
 To date, two main theoretical approaches have been pursued.
-Numerically, techniques such as exact diagonalization (ED) and density matrix renormalization group (DMRG) have been used to map out FCI phase diagrams, but these are computationally expensive and limited to small system sizes.
+Numerically, techniques such as exact diagonalization (ED) and density matrix renormalization group (DMRG) have been used to map out FCI phase diagrams [14, 15, 16], but these are computationally expensive and limited to small system sizes.
 Analytical approaches, in contrast, have primarily focused on the sufficient conditions to realize FCI phases, mostly by mimicking the lowest Landau level (LLL) physics [17].
-Notable examples include the trace condition $\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}=|\Omega_{\bm{k}}|$ restoring the Girvin-MacDonald-Platzman (GMP) algebra [15] in the small-$q$ limit [16, 17, 18, 19, 20], the ideal band condition yielding $\bm{k}$-holomorphic Bloch wavefunctions [21, 22, 23], and the notion of vortexability extending the flux attachment picture in combination of band projection [24, 25].
+Notable examples include the trace condition restoring the Girvin-MacDonald-Platzman (GMP) algebra [18] in the small-$q$ limit [19, 20, 21, 22, 23], the ideal band condition yielding $\bm{k}$-holomorphic Bloch wavefunctions [24, 25, 26], and the notion of vortexability extending the flux attachment picture in combination of band projection [27, 28].
 While insightful, these criteria apply only to narrow regions of parameter space similar to the LLL.
@@
 This motivates a central question: Can we develop a physically transparent and computationally efficient framework for diagnosing FCI stability beyond the ideal limit?
-In this Letter, we present a quantitative theory of FCI stability based on the dipole picture of composite fermions (CFs) [28, 29, 30, 31, 32, 33, 34].
+In this Letter, we present a quantitative theory of FCI stability based on the dipole picture of composite fermions (CFs) [31, 32, 33, 34, 35, 36, 37, 38].
 We construct CFs on a lattice by binding Bloch electrons to vortices residing in their own LLL, and derive an effective CF Hamiltonian using the method of preferred-charge substitution [34, 39, 40].
@@
 [37]).
-In particular, assuming a flat electron band, the trace condition further reduces this Hamiltonian to the Aharonov-Casher form [37], which guarantees the exact flatness of the lowest CF band.
-For general cases, we use the GMP expansion [32, 38], combined with our recently developed lattice projective construction of FCI states [33], to extract CF band structures.
-CF phase diagrams can then be obtained by tracking the closing of the CF band gap.
+The trace condition arises as a special case: when the electron band is flat and the trace condition is satisfied, the Hamiltonian reduces to the Aharonov-Casher (AC) form [41, 42, 43], which guarantees the exact flatness of the lowest CF band.
+To quantitatively extract the CF band structure, we go beyond the small-$q$ limit and employ the GMP expansion [35, 44], in combination with our recently developed lattice projective construction of FCI states [36].
+The collapse of the FCI phase is then signaled by the closing of the CF band gap.
 As a concrete example, we apply our theory to study FCI stability in twisted MoTe2.
@@
 As a result, the total CF charge becomes $q_{\text{CF}}=(1-c^{2})q_{e}$.
-When $\nu$ lies in the Jain sequence $\nu=p/(2ps+1)$, the CF filling factor becomes an integer $p$, and the FQHE of electrons maps to the integer quantum Hall effect of CFs.
+When $\nu$ lies in the Jain sequence $\nu=p/(2sp+1)$, the CF filling factor becomes an integer $p$, and the FQHE of electrons maps to the integer quantum Hall effect of CFs.
 This construction underlies the extended Hamiltonian theory of Murthy and Shankar, which provides a powerful tool for studying the FQHE [34].
@@
 Although we continue to call them “vortices”, they are best viewed as fictitious particles that carry fractional charge $q_{v}=-c^{2}q_{e}$, and experience an effective magnetic field set by the average Berry curvature flux, $B_{\text{eff}}=h/(eA_{\text{u.c.}})$, where $A_{\text{u.c.}}$ is the unit cell area.
-The vortex unit cell, determined by magnetic translation symmetry, is thus enlarged by a factor of $(2ps+1)$ relative to the electron unit cell.
+The vortex unit cell, determined by magnetic translation symmetry, is enlarged by a factor of $(2sp+1)$ relative to the electron unit cell.
 The CF is a bound state of an electron and a vortex that lives in the enlarged single-particle Hilbert space
@@
 […]
-$H_{0}^{P}(\bm{k},z_{v};\bm{k}^{\prime},z_{v}^{\prime})=\frac{-c^{2}}{2}\sum_{\bm{q}}V_{\bm{q}}\langle\bar{\varrho}_{e,\bm{q}}\otimes\bar{\varrho}_{v,\bm{q}}+\text{h.c.}\rangle\;,$ (7)
-where $\bar{\varrho}_{\alpha,\bm{q}}=e^{i\mathop{\mathrm{sgn}}(\alpha)\bm{q}\cdot\hat{\bm{r}}_{\alpha}}$ denotes the matrix element of the projected density operator for species $\alpha=e,v$ with $\mathop{\mathrm{sgn}}(e)=1$ and $\mathop{\mathrm{sgn}}(v)=-1$.
-Equations 5, 6 and 7 constitute the core result underlying our CF theory.
+$H_{0}^{P}$ $(\bm{k},z_{v};\bm{k}^{\prime},z_{v}^{\prime})=$
+$\frac{-c^{2}}{2}\sum_{\bm{q}}V_{\bm{q}}\langle\psi_{\bm{k}},z_{v}|\bar{\varrho}_{e,\bm{q}}\otimes\bar{\varrho}_{v,\bm{q}}+\text{h.c.}|\psi_{\bm{k}^{\prime}},z_{v}^{\prime}\rangle\;,$ (7)
+where $\bar{\varrho}_{\alpha,\bm{q}}=e^{i\mathop{\mathrm{sgn}}(\alpha)\bm{q}\cdot\hat{\bm{r}}_{\alpha}}$ denotes the projected single-particle density operator of species $\alpha=e,v$ with $\mathop{\mathrm{sgn}}(e)=1$ and $\mathop{\mathrm{sgn}}(v)=-1$.
+Equations 5, 6 and Composite Fermion Theory of Fractional Chern Insulator Stability constitute the core result underlying our CF theory.
 They describe the band structure of CFs and provide the foundation for analyzing FCI stability.
-Crucially, at filling $\nu=p/(2ps+1)$, the CFs occupy exactly $p$ bands, mapping the problem of an FCI of electrons to that of an integer Chern insulator of CFs.
+Crucially, at filling $\nu=p/(2sp+1)$, the CFs occupy exactly $p$ bands, mapping the problem of an FCI of electrons to that of an integer Chern insulator of CFs.
 This follows from the requirement that the CF periodicity must be compatible with both the electron and vortex unit cells, leading to an enlarged unit cell and hence $p$ filled CF bands.
 In the remainder of this Letter, we demonstrate that much of the CF physics is already captured by the CF single-particle Hamiltonian $H^{P}_{\text{single}}=H_{0}+H_{0}^{P}$, which encodes interaction effects through its dependence on the Coulomb potential $V_{\bm{q}}$.
-To diagonalize $H^{P}_{\text{single}}$, we can expand $\bar{\rho}_{e,\bm{q}}$ in terms of GMP density operators [32]:
-$\bar{\rho}_{e,\bm{q}}=\sum_{\bm{G}}c(\bm{q},\bm{G})\rho^{\text{GMP}}_{e,[\bm{q}]+\bm{G}}\;,$ (8)
-where $\bm{G}$ is the electron reciprocal lattice vector, $\bm{q}]$ denotes $\bm{q}$ folded back to the first Brillouin zone (BZ), and the coefficients
-$c(\bm{q},\bm{G})=\frac{1}{N_{\phi,e}}\mathop{\mathrm{Tr}}\left[\bar{\varrho}_{e,\bm{q}}\varrho^{\text{GMP}}_{e,-[\bm{q}]-\bm{G}}\right]$ (9)
-encode all lattice effects specific to the FCI.
-This so-called GMP expansion is complete and can be used for any fermionic bilinear [32, 38].
-The resulting CF band structure problem over a finite-size sample on a torus can then be solved using the lattice projective construction as described in Ref.
-[33].
-Small-$q$ Limit and the Trace Condition.—It is instructive to first consider the situation when the form factor of the density operator decays exponentially, or when $V_{\bm{q}}$ is appreciable only at small $\bm{q}$.
-For instance, the vortex density operator $\bar{\rho}_{v,\bm{q}}$ carries a form factor $e^{-(q\ell_{v})^{2}/4}$ from the LLL projection [see Eq.
-4].
-We can then expand the projected density operators for both species $\alpha=e,v$ to second order in $q$:
+Small-$q$ Limit and the Trace Condition.—It is instructive to first consider the small-$q$ expansion of the projected density operators of both species $\alpha=e,v$:
 $\bar{\varrho}_{\alpha,\bm{q}}\simeq P_{\alpha}+i\,\mathop{\mathrm{sgn}}(\alpha)\,q_{\mu}P_{\alpha}\hat{r}_{\alpha}^{\mu}P_{\alpha}-\frac{1}{2}q_{\mu}q_{\nu}P_{\alpha}\hat{r}_{\alpha}^{\mu}\hat{r}_{\alpha}^{\nu}P_{\alpha}\;.$ (8)
@@
 $\frac{\hbar^{2}}{2m_{*}\ell_{e}^{2}\ell_{v}^{2}}=\frac{c^{2}}{4}\sum_{\bm{q}}V_{\bm{q}}\bm{q}^{2}e^{-\lambda^{2}q^{2}}\;.$ (10)
-We have added a $\lambda$-parameterized exponential regulator to ensure convergence (see SM [41] for its origin).
+We have added a $\lambda$-parameterized exponential regulator to account for the vortex LLL density operator form factor and the higher-order corrections from electron density operators (see SM [47] for details).
 The terms proportional to $1/m_{*}$ in Eq.
@@
 Had we assumed the electron band carries positive Chern number $C=1$, the sign of $\Omega_{\bm{k}}$ would be reversed.
-Taken together, Eq.
-11 describes a Hofstadter problem in the $(\bm{k},z_{v})$ space: the first two terms act as a periodic scalar potential, while the last term, taking the form of an Aharonov-Casher (AC) Hamiltonian [37], describes a periodic magnetic field.
-It is clear that if the electron band is flat ($\varepsilon_{\bm{k}}=0$) and the so-called trace condition $\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}+\Omega_{\bm{k}}=0$ is satisfied, then the Hamiltonian reduces to a pure AC form, which guarantees that the lowest CF band is exactly flat.
-This perspective yields several new insights.
-First, contrary to earlier studies where $\varepsilon_{\bm{k}}$ was neglected, a nonzero electron band dispersion can help stabilize the FCI phase by compensating for deviations from the trace condition.
-Second, the trace condition is only meaningful when the lowest CF band is filled; when multiple CF bands are filled, it no longer applies as a stability criterion.
 We note that a Hamiltonian of the same form as Eq.
@@
 [37], based on fitting a postulated quadratic electron-vortex dipole model.
-Our result provides a first-principle derivation of this structure within the preferred charge formalism.
+Our result provides a first-principle derivation of this structure as the small-$q$ limit within the preferred charge formalism.
+Equation 9 describes a Hofstadter problem in the $(\bm{k},z_{v})$ space: the first two terms act as a periodic scalar potential, while the last term, taking the form of an AC Hamiltonian [41], describes a periodic magnetic field.
+It is clear that if the electron band is flat ($\varepsilon_{\bm{k}}=0$) and the trace condition $\mathop{\mathrm{tr}}\mathcal{G}_{\bm{k}}+\Omega_{\bm{k}}=0$ is satisfied, the CF Hamiltonian reduces to a pure AC form, which guarantees the lowest CF band to be exactly flat.
+However, this trace condition alone does not determine the CF band gap, whose size controls the stability of the FCI phase as we show below.
+The CF gap arises from the combined effect of three factors: (i) the scalar-potential term (including band dispersion and the trace condition), (ii) the magnetic-field term associated with Berry-curvature, and (iii) the CF effective mass, which depends on interaction details.
+CF Band Structure and CF Phase Diagram in tMoTe2—In the small-$q$ expansion Eq.
+9, we have introduced a regularization parameter for the effective mass.
+For a quantitative calculation of the CF band structure without any fitting parameter, we will work with the CF Hamiltonian in Composite Fermion Theory of Fractional Chern Insulator Stability without taking the small-$q$ limit, where the information of the Berry curvature, quantum metric and CF effective mass are all implicitly encoded in the form factor.
+To achieve this, we construct $N_{s}^{2}=(N_{e,1}\times N_{e,2})^{2}$ number of independent GMP operators $\{\rho^{\text{GMP}}_{e,\bm{q}}\}$ from the matrix elements of electronic LLL density operator to satisfy the GMP algebra [18], and expand each $\bar{\rho}_{e,\bm{q}}$ following [35, 36]
+$\bar{\rho}_{e,\bm{q}}=\sum_{\bm{G}_{\text{std}}}c(\bm{q},\bm{G}_{\text{std}})\rho^{\text{GMP}}_{e,[\bm{q}]+\bm{G}_{\text{std}}}\;.$ (11)
+Here $\sum_{\bm{G}_{\text{std}}}$ runs over $N_{s}$ number of reciprocal vectors, and $\bm{q}]$ denotes $\bm{q}$ folded back to the first Brillouin zone (BZ).
+The coefficients (see SM [47] for derivation)
+$c(\bm{q},\bm{G}_{\text{std}})=\frac{1}{N_{s}}\mathop{\mathrm{Tr}}\left[\bar{\varrho}_{e,\bm{q}}\varrho^{\text{GMP}}_{e,-[\bm{q}]-\bm{G}_{\text{std}}}\right]$ (12)
+encodes all lattice effects specific to the FCI.
+This so-called GMP expansion is complete and can be used for any fermionic bilinear [35, 44].
+The resulting CF band structure problem can then be solved by diagonalizing the full-$q$ CF Hamiltonian.
 Figure 1: CF phase diagram (a), $6\times 4$ ED phase diagram (b), geometric fluctuations (c), and typical CF bands in twisted MoTe2 (d–f).
 The color maps in (a) and (b) represent the CF band gap and the many-body gap obtained from ED, respectively.
-Panel (c) shows the standard deviation of the effective magnetic field $\Omega_{\bm{k}}$ and the effective scalar potential of the CF Hamiltonian (11).
-The standard deviation of the trace condition is also shown.
+Panel (c) shows the dimensionless standard deviation of the effective magnetic field $\Omega_{\bm{k}}$, traditional trace condition, and the effective scalar potential of the CF Hamiltonian (9) at $\epsilon=15$, normalized by the moiré unit cell volume $A_{\text{u.c.}}$ and the effective CF mass $m_{*}\approx 4.87$ meV measured from the calculated CF spectrum close to the ideal limit $\theta=3.55^{\circ}$.
 Panels (d)–(f) show the two lowest CF bands for three representative points labeled in (a), computed on a $9\times 9$ electronic sample, with colors represents the band widths.
@@
 The extra threefold degeneracy within CF’s reduced BZ, known as the symmetry fractionalization pattern due to the interplay between topological orders and magnetic translation symmetries [48, 49, 50, 36], is also explicitly visible in our calculation.
-CF Band Structure and CF Phase Diagram in tMoTe2.—We now apply our theory to the continuum model of the $R$-stacked twisted MoTe2 bilayer.
+We now apply our theory to the continuum model of the $R$-stacked twisted MoTe2 bilayer.
 We assume the system is valley polarized, and focus on the effective Hamiltonian at the $K$ valley, given by [51]:
@@
 The momenta $\bm{K}_{b}=\frac{1}{3}(\bm{g}_{4}+\bm{g}_{5})$ and $\bm{K}_{t}=\frac{1}{3}(\bm{g}_{3}+\bm{g}_{4})$ are the $K$ points of the top/bottom layer and $\theta$ is the twist angle.
-All parameters are from Ref.
-[46].
-The Coulomb potential with dual-gate screening is taken to be $V_{\bm{q}}=\frac{1}{A_{\text{u.c.}}}\frac{e^{2}\tanh(|\bm{q}|d)}{2\epsilon\epsilon_{0}|\bm{q}|}$, where $d$ is the gate-to-sample distance.
+All parameters follows Ref.
+[52]: $m_{*}=0.6m_{e}$, $a=3.52$ Å, $v=20.8~\text{meV}$, $\psi=107.7^{\circ}$, $w=-23.8~\text{meV}$.
+The Coulomb potential with dual-gate screening is taken to be $V_{\bm{q}}=\frac{1}{A_{\text{u.c.}}}\frac{e^{2}\tanh(|\bm{q}|d)}{2\epsilon\epsilon_{0}|\bm{q}|}$, where $d=300$ Åis the gate-to-sample distance.
 Let us focus on the hole filling $\nu=-2/3$ of the topmost valence band.
 We take a particle-hole transformation to convert it to an equivalent electron-doped model at filling $\nu=1/3$, squeezing out a Hartree-term $\varepsilon_{\bm{k}}^{\text{Hartree}}\equiv-\sum_{\bm{G}}V_{\bm{G}}(\sum_{\bm{k}^{\prime}}\bar{\varrho}_{-\bm{G}}([\bm{k}^{\prime}]))\bar{\varrho}_{\bm{G}}([\bm{k}])$ that is absorbed into the original band dispersion $\varepsilon_{\bm{k}}\mapsto-\varepsilon_{\bm{k}}-\varepsilon_{\bm{k}}^{\text{Hartree}}$.
-We construct the CF Hamiltonian using the preferred charge substitution (3), and solve it via the GMP expansion (8) and our recently developed lattice projective construction [33].
+We construct the CF Hamiltonian using the preferred charge substitution (3), and solve it via the GMP expansion (11) and our recently developed lattice projective construction [36] without resorting to the small-$q$ limit.
 Figure 1(d) shows the CF band structure at $\theta=3.55^{\circ}$ and $\epsilon=15$.
@@
 1(a).
-The resulting CF phase diagram shows remarkable agreement with ED results in Fig.
-1(b), capturing not only the shape of the FCI region but also the location of the maximal many-body gap.
+The resulting CF phase diagram shows close agreement with ED results in Fig.
+1(b), capturing not only the shape of the FCI region but also the location of the maximal many-body gap  [^1] Note that when comparing the CF phase diagram with the ED phase diagram, one should keep in mind that the CF gap corresponds to a charge excitation but the ED gap measures the magnetoroton minimum [18]..
 This agreement strongly supports the predictive power of our CF-based approach and confirms that the CF band gap serves as a reliable proxy for FCI stability.
@@
 As $\epsilon$ increases, the inverse CF mass $1/m_{*}$ decreases [see Eq.
-12], leading to two key effects illustrated in Fig.
+10], leading to two key effects illustrated in Fig.
 1(e).
@@
 1(c), both the scalar potential [the first two terms in Eq.
-11] and the effective magnetic field $\Omega_{\bm{k}}$ exhibit increasing fluctuations with larger $\theta$.
+9] and the effective magnetic field $\Omega_{\bm{k}}$ exhibit increasing fluctuations with larger $\theta$.
 As a result, CF bands develop significant dispersion, as shown in Fig.
@@
 The agreement is remarkable: within the FCI phase, the projected states achieve overlaps exceeding 90%, with a peak value of $(99.37\pm 0.06)\%$ at $\theta=3.5^{\circ}$, where the many-body gap reach its maximum.
-As $\theta$ increases, the energy of eigenstates outside the topological manifold descend and eventually become the new ground states, signaling the breakdown of the FCI phase [Fig.
+As $\theta$ increases, the eigenstates outside the topological manifold descend in energy and eventually become the new ground states, signaling the breakdown of the FCI phase [Fig.
 2(c)].
-We continue to track the states at the three center-of-mass momenta where FCI’s topological ground states reside under this transition, and find that their overlaps with the projected electronic states remain finite.
-In contrast, the overlaps with the new ground states are negligible.
+However, we continue to track the ED eigenstates at the three original momenta and find that their overlaps with the projected CF states remain substantial even after they are no longer the ground states.
+In contrast, the new ground states have negligible overlap with the three projected CF states.
 This result confirms that our framework not only qualitatively reproduces the FCI phase diagram, but also provides quantitatively accurate many-body wavefunctions through the electronic projection defined in Eq.
@@
 Another promising direction is to explore whether residual CF interactions can drive symmetry-breaking phases, such as CF charge orders or CF pairings, beyond the simple band picture.
-Additionally, incorporating band-mixing effects remains an important task.
+Additionally, generalizing our framework to composite fermi liquid (CFL) states in FCI [59, 60], and incorporating band-mixing effects [61] remain to be important tasks.
 These extensions could further deepen our understanding of FCI physics and strengthen the connection between microscopic theory and experimental observations.
@@
 DE-SC0019443.
+This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S.
+Department of Energy under Contract No.
+DE-AC02-05CH11231 using NERSC award BES-ERCAP0033507.